OpenStudy (anonymous):

Use the “difference of squares” rule to factor the following expression: 16-y^2

OpenStudy (saifoo.khan):

-(-16+y^2) -(y^2 - 16) -(y-4)(y+4)

OpenStudy (anonymous):

A^2 - b^2 = (a+b)(a-b). In this case, 16 - y^2 = (4+y)(4-y)

OpenStudy (unklerhaukus):

\[ (a+b)(a-b)=a^2-b^2\]

OpenStudy (saifoo.khan):

It's same thing Kumar.

OpenStudy (anonymous):

I am not blind Saifoo

OpenStudy (saifoo.khan):

Lol, i never said that you are blind.

OpenStudy (anonymous):

Well you said it's same thing liek a smartass

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

Please hit Good ANswer @asker.

OpenStudy (anonymous):

Good answer bhahahahaha nice joke

OpenStudy (saifoo.khan):

Fail Joke. :p

OpenStudy (anonymous):

16-y^2 = 4^2 - y^2 = ( 4 - y) ( 4 + y)